Calculate the molality of a solution containing 10.8g of ethylene glycol (C2H6O2) in 360g of water. 


1
Expert's answer
2021-02-25T03:20:09-0500

n(C2H6O2)=m/Mr=10.8/62=0.17 mol

molality=n(C2H6O2)/m(H2O(kg))= 0.17/0.36=0.48 mol/kg


Answer: the molality of solution is 0.48 mol/kg